A shipment of 8 similar microcomputers to a retail outlet contains 3 defectives. If a school makes a random purchase of 2 of these computers, nd the probability distribution for the number of defectives. also obtain the cumulative probability distribution.


1
Expert's answer
2021-01-05T14:27:28-0500

"p = \frac{3}{8} = 0.375"

n = 2

X ~ Bin(n, p)

X takes value 0, 1, 2

P(X=x) = nCxpx(1-p)n-x

P(X=0) = 2C00.3750(1-0.375)2-0

"= \frac{2!}{0!2!}(0.375)^0(0.625)^2 \\

= 0.390625"

P(X=1) = 2C10.3751(1-0.375)2-1

"= \frac{2!}{1!1!}(0.375)^1(0.625)^1 \\

= 0.468750"

P(X=2) = 2C20.3752(1-0.375)2-2

"= \frac{2!}{2!0!}(0.375)^2(0.625)^0 \\

= 0.140625"

Probability distribution of X:



Fx(0) = 0.390625

Fx(1) = P(X=0) + P(X=1) = 0.390625 + 0.468750 = 0.859375

Fx(2) = Fx(1) + P(X=2) = 0.859375 + 0.140625 = 1

Cumulative probability distribution: